OpenStudy (anonymous):

Solve correct to four decimal places: 3^x =4^(x-7) What do I do since the bases are not the same?

OpenStudy (anonymous):

Take the log of both sides.

OpenStudy (anonymous):

a raft of algebra

OpenStudy (anonymous):

start with \[\ln(3)x=(x-7)\ln(4)\] and then solve for \(x\)

OpenStudy (anonymous):

don't be confused, \(\ln(3)\) and \(\ln(4)\) are just some constants (numbers)

OpenStudy (anonymous):

you got the algebra?

OpenStudy (anonymous):

Would I divide log(3) from both sides?

OpenStudy (anonymous):

I canceled the logs on both sides and ended up with 3x= 4x-14 x=14